you need rm20 000 for your education by the end of 4 years. you have rm5 000 and wish to save balance by deposits every year. calculate the amount that you have to save at the end of each year with 8 percent interest.


1
Expert's answer
2022-03-04T02:51:02-0500

Amount remaining= 20000-5000= 15,000

FV= 15,000

"PV=\frac{FV}{(1+0.08)^4}"

"=\frac{15000}{(1+0.08)^4}"

"=\frac{15000}{1.36048896}= 11,025.45"